OpenStudy (anonymous):

rt(3) tan 3θ + 1 = 0

OpenStudy (anonymous):

\[\sqrt{3} \tan 3\theta +1 = 0\]

OpenStudy (mertsj):

\[\tan 3\theta=\frac{-\sqrt{3}}{3}\] \[3\theta=\frac{5 pi}{6},\frac{11 pi }{6}, \frac{17 pi}{6}, \frac{23pi}{6}, \frac{29pi}{6}, \frac{35pi}{6}\] \[\theta=\frac{5pi}{18}, \frac{11pi}{18}, \frac{17pi}{18}, \frac{23pi}{18}, \frac{29pi}{18}, \frac{23pi}{18} + 2pin\]

OpenStudy (mertsj):

Whoops. Forgot 35 pi/18

OpenStudy (anonymous):

tg(3.θ)=−3√/3→3.θ=(K).pi−pi/6 k is interger

OpenStudy (anonymous):

\[k pi/3+ pi/18\] is what i entered and its wrong? Any idea why?

OpenStudy (anonymous):

-pi/18

OpenStudy (anonymous):

see?

OpenStudy (anonymous):

yepp thankss!